OpenStudy (anonymous):

f(x) = ax^3+bx^2+cx+d find a,b,c,d so that the function has a local maximum at f(-1) = 2 and has a local minimum at f(1) = -1 HARDEST QUESTION EVER I CAN NOT FIGURE THIS OUT!!! PlZ I NEED THE BEST OF THE BEST

OpenStudy (anonymous):

And I tell u this the simplest of all I have ever seen.

OpenStudy (anonymous):

Really?!

OpenStudy (anonymous):

I think i just don't know what it's asking

OpenStudy (anonymous):

Ok here is how u solve it f(-1) = 2 =-a+b-c+d f(1) = -1 = a+b+c+d And since the above two points are local minimum and maximums, this must hold true f'(-1) = 0 and f'(1) = 0 But f'(x) = 3ax^2+2bx+c Thus f'(-1) = 0=3a-2b+c f'(1) = 0 = 3a+2b+c Now you have four equations with four variable. Will u finish from this?

OpenStudy (anonymous):

oh yes, thank you

OpenStudy (anonymous):

your a genious

OpenStudy (anonymous):

you deserve a metal